Perlor, a musical instruments manufacturer in Indonesia, make and sell guitars. Let’s assume they operate at maximum capacity, making guitars seven days a week, 52 weeks per year, with a production result of 100 guitars per week. Perlor ship all guitars they can produce as soon as they can, keeping no safety stock.

The guitars are shipped to Germany, where a wholesaler distributes them to all over Europe. Sales in units is the same as the production in Indonesia.

The value of a guitar leaving production is €200. The cost of capital is estimated to 15%. The cost for holding inventory is 10%.

Shipments dispatch every eight weeks. The shipping time is six weeks, and the cost is €5/unit.

In Germany, the wholesaler keeps a safety stock corresponding to a week of sales.

6a. What is the average inventory in Indonesia?

  units  

6b. What is the average inventory during shipping?  units

6c. What is the average inventory in Germany?  units  

6d. What is the inventory holding cost in Indonesia?  €/year  

6e. What is the inventory holding cost during transportation?  €/year  

6f. What is the inventory holding cost in Germany?  €/year  

6g. What is the shipping cost?  €/year

6h. What is the total cost , considering the total inventory carrying cost and shipping?  €/year

Please answer all 8 Parts. correct and exact answer . i have to submit answer online
Answer #1

Ans. 6(a). Opening Inventory = 0 Units and Closing Inventory = (4 x 100) = 400 Units

Therefore, Average inventory in Indonesia = 400/2 = 200 Units

Ans. 6(b). Average inventory during shipping = 400/2 = 200 Units

Ans. 6(c). Average inventory in Germany = 100 Units

Ans. 6(d). Inventory holding cost in Indonesia = (200 x 5200 x 10%) x 8/52 = 16000

Ans. 6(e). Inventory holding cost during transportation = (200 x 5200 x 10%) x 6/52 = 12000

Ans. 6(f). Inventory holding cost in Germany = (5 x 5200) x 6/52 = 3000

Ans. 6(g). Shipping cost = 5 x 5200 = 26000

Ans. 6(h). Total Cost = 16000 + 3000 + 26000 = 45000